# Time-Varying Factor-Augmented Models for Volatility Forecasting

### v2

@@ 4.3.2 Economic Implication
$\widehat{\mathrm{UoW}}=\frac{1}{T}\sum_{t=1}^{T}\!\left[\frac{SR^{2}}{\gamma}\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}-\frac{SR^{2}}{2\gamma}\!\left(\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}\right)^{2}\right],$ (16)
where we follow bollerslev2018risk in setting the target Sharpe ratio to $\mathrm{SR}=0.4$, a typical long-run estimate for volatility-timing strategies, and choose risk aversion $\gamma=2$ to represent moderate investor preferences. Higher $\mathrm{UoW}$ denotes greater expected utility for a volatility-timing investor.
4.3.3 Statistical Significance
We assess systematic improvements with the Diebold–Mariano testdiebold2002comparing . For each horizon $h$, we compute the one‐period loss differential:
$d_{t}\;=\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{aug}}\bigr)\;-\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{RW}}\bigr),$ (17)
@@ 4.3.3 Statistical Significance
$\mathrm{DM}=\frac{\bar{d}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d})}},\quad\bar{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t},$ (19)
@@
5 Experiment Result
Table 1: Cryptocurrency Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
@@ 5 Experiment Result
Note: All MSE values are in units of $10^{-4}$.   Improvement:  $\geq+15\%$  $\geq+10\%$  $\geq+5\%$  $0\%$  $\leq-5\%$
5.1 A Universal Performance Boost Across All Models and Horizons
@@
5.2 Value Across Markets: Robust in Equities, Super‑Charged in Crypto
@@
5.3 Statistically and Economically Significant Gains
Factor augmentation delivers gains that are both substantial and statistically significant in both predictive accuracy and economic value. We demonstrate this on the crypto 7-day horizon task, where our framework shows the largest improvements. Three unaugmented baselines are employed: the Random Walk, an "Individual" model with asset-specific coefficients, and an "Panel" model that share coefficients across assets.
@@ 6.1 Cointegration Test
Table 2: Cointegrating Vector Coefficients for Crypto
@@
6.2 Volatility‑Based Pairs Trading
@@
$S_{t}=\ln P_{\text{ADA},t}-\beta\,\ln P_{\text{ETH},t},$ (20)
@@
6.2.1 Trading Signal Generation
Trading signals are generated from a rolling z-score of the spread, where $Z_{t}=\frac{S_{t}-\mu_{S_{t}}}{\sigma_{S_{t}}}$. We initiate a long position when $Z_{t}$ < -1.5 and a short position when $Z_{t}$ > +1.5, closing the position when the z-score reverts to zero. To limit turnover, each trade is held for at least one day. The rolling window length ($W$) is set to 70 days for the statistical models to balance responsiveness to regime changes with statistical stability, while a shorter window of 30 days is used for the LSTM to align with its input structure.
6.2.2 Volatility-based Position Sizing
@@
$\widehat{\sigma}_{\text{spread},t}\;=\;\sqrt{252}\cdot\sqrt{\widehat{\sigma}_{\text{ADA},t}^{2}+\beta^{2}\widehat{\sigma}_{\mathrm{ETH},t}^{2}-2\beta\,\widehat{\mathrm{Cov}}(r_{\text{ADA},t},r_{\text{ETH},t})}.$ (21)
@@ 6.3 Backtest Results
Table 3: Out-of-Sample Backtest for ADA-ETH Pair
@@
6.3.1 Dual Gains: Amplifying Return and Sharpening Risk
@@
6.3.2 Robustness in All Market Conditions: Enhancing Gains, Reversing Losses.
@@
7 Conclusion
This paper proposes a novel, model-agnostic Factor-Augmented Volatility Forecast framework, resolving the trade-off between computationally infeasible multivariate models, overly simplistic univariate models, and rigid static-factor models. By extracting a compact set of dynamic factors and their time-varying loadings directly from realized volatilities, our framework allows any baseline model to efficiently capture evolving market-wide co-movements. The framework’s robust performance—evidenced by substantial gains in forecast accuracy and superior risk-adjusted returns—combined with its computational efficiency and interpretability, establishes it as a valuable and practical solution for dynamic risk management.
Appendix
8 Equity Market: Forecast Performance
Table 4: Equity Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
@@ 8 Equity Market: Forecast Performance
Note: All MSE values are in units of $10^{-4}$.   Improvement:  $\geq+15\%$  $\geq+10\%$  $\geq+5\%$  $0\%$  $\leq-5\%$
9 Statistical Significance
(a) Out-of-Sample $R^{2}$ and MSE-based DM Test for 7-day Crypto

### v3

@@ 4.3.2. Economic Implication
(16) $\widehat{\mathrm{UoW}}=\frac{1}{T}\sum_{t=1}^{T}\!\left[\frac{SR^{2}}{\gamma}\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}-\frac{SR^{2}}{2\gamma}\!\left(\frac{RV^{7}_{t+7}}{\widehat{RV}^{7}_{t+7}}\right)^{2}\right],$
where we follow (Bollerslev et al., 2018) in setting the target Sharpe ratio to $\mathrm{SR}=0.4$, a typical long-run estimate for volatility-timing strategies, and choose risk aversion $\gamma=2$ to represent moderate investor preferences. Higher $\mathrm{UoW}$ denotes greater expected utility for a volatility-timing investor.
4.3.3. Statistical Significance
We assess systematic improvements with the Diebold–Mariano test(Diebold and Mariano, 2002). For each horizon $h$, we compute the one‐period loss differential:
(17) $d_{t}\;=\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{aug}}\bigr)\;-\;L\bigl(RV^{h}_{t+h},\widehat{RV}^{h}_{t+h,\mathrm{Baseline}}\bigr),$
@@ 4.3.3. Statistical Significance
(19) $\mathrm{DM}=\frac{\bar{d}}{\sqrt{\widehat{\mathrm{Var}}(\bar{d})}},\quad\bar{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t},$
@@
5. Experiment Result
Table 1. Cryptocurrency Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
@@ 5. Experiment Result
Note: All MSE values are reported in units of $10^{-4}$.  Improvement:   $\geq+15\%$     $\geq+10\%$    $\geq+5\%$     > $0\%$     $\leq-5\%$
5.1. A Universal Performance Boost Across All Models and Horizons
@@
5.2. Value Across Markets: Robust in Equities, Super‑Charged in Crypto
@@
5.3. Statistically and Economically Significant Gains
Factor augmentation delivers gains that are both substantial and statistically significant in both predictive accuracy and economic value. We demonstrate this on the crypto 7-day horizon task, where our framework shows the largest improvements. Three unaugmented baselines are employed: the Random Walk, an ”Individual” model with asset-specific coefficients, and an ”Panel” model that share coefficients across assets.
@@ 6.1. Cointegration Test
Table 2. Cointegrating Vector Coefficients for Crypto
@@
6.2. Volatility‑Based Pairs Trading
@@
(20) $S_{t}=\ln P_{\text{ADA},t}-\beta\,\ln P_{\text{ETH},t},$
@@
6.2.1. Trading Signal Generation
Trading signals are generated from a rolling z-score of the spread, where $Z_{t}=\frac{S_{t}-\mu_{S_{t}}}{\sigma_{S_{t}}}$. We initiate a long position when $Z_{t}$ ¡ -1.5 and a short position when $Z_{t}$ ¿ +1.5, closing the position when the z-score reverts to zero. To limit turnover, each trade is held for at least one day. The rolling window length ($W$) is set to 70 days for the statistical models to balance responsiveness to regime changes with statistical stability, while a shorter window of 30 days is used for the LSTM to align with its input structure.
6.2.2. Volatility-based Position Sizing
@@
(21) $\widehat{\sigma}_{\text{spread},t}\;=\;\sqrt{252}\cdot\sqrt{\widehat{\sigma}_{\text{ADA},t}^{2}+\beta^{2}\widehat{\sigma}_{\mathrm{ETH},t}^{2}-2\beta\,\widehat{\mathrm{Cov}}(r_{\text{ADA},t},r_{\text{ETH},t})}.$
@@ 6.3. Backtest Results
Table 3. Out-of-Sample Backtest for ADA-ETH Pair
@@
6.3.1. Dual Gains: Amplifying Return and Sharpening Risk
@@
6.3.2. Robustness in All Market Conditions: Enhancing Gains, Reversing Losses.
@@
7. Conclusion
This paper proposes a novel, model-agnostic Factor-Augmented Volatility Forecast framework, resolving the trade-off between computationally infeasible multivariate models, overly simplistic univariate models, and rigid static-factor models. By extracting a compact set of dynamic factors and their time-varying loadings directly from realized volatilities, our frasmework allows any baseline model to efficiently capture evolving market-wide co-movements. The framework’s robust performance—evidenced by substantial gains in forecast accuracy and superior risk-adjusted returns—combined with its computational efficiency and interpretability, establishes it as a valuable and practical solution for dynamic risk management.
Acknowledgements.
Elynn Chen’s research is supported in part by the NSF Award 2412577.
Appendix A Equity Market: Forecast Performance
Table 4. Equity Volatility Forecast Accuracy: 1- and 7-Day Out-of-Sample Results
@@ Appendix A Equity Market: Forecast Performance
Note: All MSE values are reported in units of $10^{-4}$.  Improvement:   $\geq+15\%$     $\geq+10\%$    $\geq+5\%$     > $0\%$     $\leq-5\%$
Appendix B Statistical Significance
(a) Out-of-Sample $R^{2}$ and MSE-based DM Test for 7-day Crypto.